A piece of wire 20 m long is cut into two pieces. One piece is bent into a square and the other is bent into an equilateral triangle. How should the wire be cut so that the total area enclosed is: (a) maximum? (b) minimum?Β 


1
Expert's answer
2021-08-11T19:38:47-0400

i)

Let x meters be the side of square , so 4"x" meters be the total amount of wires used for square .(with 0"\leq4x\leq20 \implies 0\leq x\leq5)." For equilateral triangle 20-4"x" , wire remains .and the side will be "\dfrac{20-4x}{3}"


Given the side l of an equilateral triangle , the height is , is using the pitagoras theorem , we get -"h=\dfrac{l{\sqrt3}}{2}," so the area of triangle will become ,


"A=\dfrac{l}{2}\dfrac{l{\sqrt 3}}{2}\dfrac{1}{2}"


so the total area is -


"A=x^{2}+(\dfrac{20-4x}{3})^{2}\dfrac{\sqrt 3}{4}"


"A{'}=2x+\dfrac{\sqrt 3}{4}\dfrac{1}{9}.2(20-4x).(-4)=2x-\dfrac{2{\sqrt 3}}{9}(20-4x)"


"A^{'}>0 \ if"

"=2x-\dfrac{2{\sqrt3}}{9}(20-4x)>0"


"x>\dfrac{40\sqrt 3}{2(9+4{\sqrt 3})}"


"x>\dfrac{20(3{\sqrt 3}-4)}{11}="

So the function area is decreasing in [0,"x)" and growing in (x,5] so the point is "(\bar x,5]" is minimum of the function . The conclusion is that maximum (absolute maximum) is minimum of side of the function of the interval of function "x" ,0,5 .


if x=0 , then the area will be - "A=19.2"

if x=0 , then the area will be "x" =25.

in case the maximum area will be there where all the wire used as square .


ii)

Let the two pieces be "x\ and\ y."

Given,

"=x+y=20"


Perimeter of square ="x"

x

side"\ =\dfrac{x}{4}"


Area of square ="\dfrac{x^{2}}{16}"


Perimeter of triangle "=y"


Side"\ =\dfrac{y}{3}"


We know , that area of triangle is given by "=\dfrac{\sqrt3}{4}a^{2}"


"=\dfrac{\sqrt3}{4}(\dfrac{y}{3})^{2}=\dfrac{\sqrt{3}}{36}y^{2}"


"z=" area of square +area of triangle


"z=\dfrac{x^{2}}{16}+\dfrac{\sqrt{3}}{36}y^{2}"


Now replacing the value of "y,"


Replacing "y=20-x,"


"z=\dfrac{x^{2}}{16}+\dfrac{\sqrt3}{36}(20-x)^{2}"


"=\dfrac{dz}{dx}=\dfrac{x}{8}-\dfrac{\sqrt3}{18}(20-x)"


equate ,"\dfrac{dz}{dx}=0"


we have ,


"=\dfrac{x}{8}-\dfrac{\sqrt 3(20-x)}{18}=0"


"=\dfrac{x}{8}=\dfrac{\sqrt 3(20-x)}{18}"


Solving the above equation , we get -


"x=\dfrac{80\sqrt 3}{9+4\sqrt 3}"


"y=20-\dfrac{80\sqrt 3}{9+4\sqrt 3}"


"=\dfrac{d^{2}z}{dx^{2}}=\dfrac{1}{8}+\dfrac{\sqrt 3}{18}>0"



Thus , z is minimum when "x=" "\dfrac{80\sqrt 3}{9+4\sqrt 3}" "y=\dfrac{180}{9+4\sqrt 3}" .